OpenStudy (anonymous):

HELP PLEASE! (note: the ~ symbol is representing a square root sign) Help please. (need to simplify it and assume all variables are real numbers) ^3~27(x+3)^3= I came up with: 9x^2+3~3+9~3(x+3)

OpenStudy (anonymous):

@Hoa did I do it correctly?

OpenStudy (anonymous):

Satellite! thank goodness... HELP PLEASE?

OpenStudy (anonymous):

\[\sqrt[3]{27(x+3)^3}=3(x+3)\]

OpenStudy (anonymous):

what did I do wrong? :(

OpenStudy (anonymous):

the three is the "index" it means the cubed root

OpenStudy (anonymous):

ohhhhh.. *slaps forehead*

OpenStudy (anonymous):

the cubed root of 27 is 3, because \(3^3=27\)

OpenStudy (anonymous):

and the cubed root of \((x+3)^3\) is \(x+3\) because... because \((x+3)^3=(x+3)^3\)

OpenStudy (anonymous):

Thanks for EXPLAINING. :)